OpenStudy (anonymous):

Write the equation of the given circle. center (3, 3) radius of 2

OpenStudy (anonymous):

The general equation of a circle is: \[(x-h)^2+(y-k)^2=r^2\] Where (h,k) is the center and r is the radius

OpenStudy (anonymous):

Are you able to plug those values in?

OpenStudy (anonymous):

(x-3)^2+(y-3)^2=12

OpenStudy (anonymous):

Check your radius

OpenStudy (anonymous):

=4

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Thanks very helpful

OpenStudy (anonymous):

You're welcome!